# How Good is Your Wikipedia?

### v1

@@ 5.1 Monolingual Pretraining
To isolate the impact of data quality on downstream performance, we pretrain monolingual models, with a scaled down DeBERTa He et al. (2021) architecture on 51 Wikipedias from the 4 tiers. These languages are selected based on the availability of downstream evaluation tasks. Training small monolingual models minimises the risk of confounding factors affecting the results, while also being very efficient to train.
We study the effect of data pruning by pretraining models on three versions of Wikipedia: (1) the entire Wikipedia of the language without any filtering (raw_wiki), (2) filtered Wikipedia using the primary filters (+primary), and (3) further pruned using the heuristic filters (+heuristic), as described in §3. The full experimental pipeline is illustrated in Figure 5.
To understand the saliency of each class of heuristic filters detailed in §3.2, we sort the articles based on their total score and split them into high and low quality partitions such that both partitions have a roughly equal number of characters. We then pretrain monolingual models over these partitions for each class. This ensures that any differences in performance can be attributed to data quality and not size.
For every language, we pretrain 9 monolingual models. All models are subsequently fine-tuned on MasakhaNER 2.0 Adelani et al. (2022), MasakhaPOS Dione et al. (2023), MasakhaNEWS Adelani et al. (2023), AfriSenti-Twitter Muhammad et al. (2023) and SIB200 Adelani et al. (2024), based on availability for every language. Details about model and tokeniser configurations and the entire training regime are in Appendix A.
5.1.1 Effect of Data Quality Filtering
Comparing the effect of data quality filtering across various downstream tasks reveals that primary filtering generally is not detrimental to accuracy. Specific results per language and per task may vary. MasakhaNER (Table 2) on average has similar performance for no filtering and primary filtering, but has a slight performance drop with heuristic filtering. MasakhaNEWS (Table 3), on the other hand, has a stronger preference for primary filtering. AfriSenti-Twitter (Table 4) has the same average performance with or without any filtering.
We note that these averages are not entirely representative of trends with individual Wikipedias. For example with MasakhaNEWS, while there are minimal differences across the partitions for most languages, om and ti get a substantial performance boost from primary filtering. Conversely, rn performance after heuristic filtering is much worse than no filtering at all.
For AfriSenti-Twitter, although the average scores across partitions are the same, there is never a performance decline with +primary, except for ary. In fact, am has a 6% performance boost from no filtering to heuristic filtering. The ary performance drop can be explained by the fact that the ary subset in AfriSenti-Twitter contains script switching between Arabic and Latin. None of the models, except raw_wiki, is able to tokenise Latin characters due to the script filtering step which removes all Latin characters from the vocabulary. This connects to the discussion in §2.3 about what data quality actually means (more in §6).
@@ 5.1.1 Effect of Data Quality Filtering
Lang raw_wiki +primary +heuristic
pcm 0.77±0.00 0.76±0.00 0.76±0.00
ha 0.82±0.00 0.81±0.00 0.81±0.00
ig 0.85±0.00 0.85±0.00 0.84±0.00
sn 0.92±0.00 0.92±0.00 0.92±0.00
rw 0.76±0.05 0.79±0.01 0.77±0.05
tw 0.76±0.00 0.76±0.01 0.74±0.00
wo 0.75±0.01 0.73±0.01 0.72±0.01
ny 0.87±0.00 0.87±0.00 0.86±0.00
ee 0.81±0.00 0.81±0.01 0.79±0.00
bm 0.74±0.01 0.74±0.01 0.72±0.01
fon 0.74±0.01 0.74±0.00 0.76±0.01
sw 0.90±0.00 0.90±0.00 0.90±0.00
zu 0.83±0.01 0.82±0.01 0.81±0.01
lg 0.84±0.00 0.85±0.00 0.83±0.00
tn 0.80±0.01 0.80±0.00 0.79±0.01
yo 0.84±0.00 0.84±0.00 0.83±0.00
Avg 0.81 0.81 0.80
@@
Lang raw_wiki +primary +heuristic
pcm 0.90±0.04 0.91±0.00 0.87±0.04
ha 0.88±0.01 0.86±0.01 0.86±0.01
ig 0.85±0.01 0.85±0.01 0.84±0.01
sn 0.89±0.01 0.90±0.01 0.89±0.01
so 0.71±0.01 0.71±0.01 0.70±0.01
ln 0.78±0.01 0.76±0.01 0.76±0.01
om 0.67±0.30 0.82±0.01 0.81±0.30
ti 0.60±0.04 0.65±0.02 0.62±0.04
sw 0.81±0.01 0.82±0.00 0.82±0.01
am 0.84±0.04 0.87±0.01 0.85±0.04
lg 0.86±0.01 0.86±0.02 0.84±0.01
rn 0.75±0.01 0.76±0.02 0.71±0.01
yo 0.84±0.03 0.86±0.01 0.86±0.03
Avg 0.80 0.82 0.80
@@
Lang raw_wiki +primary +heuristic
pcm 0.61±0.01 0.61±0.01 0.61±0.01
ha 0.72±0.01 0.72±0.01 0.73±0.01
ig 0.74±0.00 0.75±0.00 0.75±0.00
rw 0.60±0.04 0.61±0.01 0.59±0.04
tw 0.62±0.01 0.64±0.01 0.64±0.01
ts 0.48±0.04 0.49±0.03 0.50±0.04
sw 0.55±0.03 0.59±0.02 0.56±0.03
am 0.40±0.07 0.41±0.04 0.46±0.07
ary 0.46±0.01 0.40±0.01 0.40±0.01
yo 0.68±0.01 0.68±0.01 0.68±0.01
Avg 0.59 0.59 0.59
@@
With SIB200 we compare the effectiveness of data quality filtering across the tiers defined in §4. Table 5 contains the results grouped by tier. On average, languages in Tier 1 have similar performance with raw_wiki and +primary, but experience a slight drop with +heuristic, a trend also seen in MasakhaNER and MasakhaNEWS. More languages in Tier 2 have the best performance with +heuristic, even though on average results are similar across partitions. On average, Tier 3 has better performance with raw_wiki. Notably, lij and ace show a 7% and 10% performance drop with heuristic filtering respectively, impacting the overall average. However, more Wikipedias achieve their highest scores with filtering than when trained on raw_wiki (5 with +primary, 5 with +heuristic and 4 with raw_wiki). For the Tier 4 Wikipedias, while the average does not demonstrate a strong preference for one partition over another, individual languages differ a lot. Cebuano, for instance, benefits substantially from primary filtering, with a 12% performance gain. It can be attributed to the deduplication step that removes most bot generated articles.
@@
Lang raw_wiki +primary +heuristic
@@
5.1.2 Saliency of Heuristic Filters
Results for models trained on low and high quality partitions from each class of heuristic filters are illustrated in Figure 6. Scores for each model and task are in Appendix B. Notably, we find that models pretrained on essentially half the data of the +primary filtering and raw_wiki models perform just as well for most tasks. Recall that with models in Tables 2, 3, 4 and 5 performance with filtering is on par with performance without. This indicates that we do not need large amounts of data to achieve competitive performance on downstream tasks, with quality mattering more than quantity during pretraining.
Comparing the performance difference between each heuristic filter, there is a signal from the distribution of the pretraining data to downstream performance, but it is not consistent. Thus, we cannot draw strong conclusions about the efficacy of each individual heuristic filter. There is little variance across models, across tasks, which can also explain why the heuristic filtering step also has little effect on performance.
Figure 6: Fine-tuning results for models trained on the absolute_low (blue), absolute_high (orange), relative_low (green), relative_high (red), entropy_low (purple), and entropy_high (brown) partitions.
5.2 Language adaptation
While the tiny models allow us to fully control the experimental variables, we also test the effect of our filtering methods in a more practical scenario. Language adaption of existing multilingual models to an unseen language is a common practice for low-resource languages Pfeiffer et al. (2020, 2022). To test this scenario, we follow the experimental protocol from Alabi et al. (2022) in training AfroXLM-R and test the efficacy of filtering Wikipedia by performing language adaption on XLM-R base Conneau et al. (2020) on unseen languages from our experimental cohort. We use the original XLM-R tokeniser and further pretrain XLM-R on Wikipedia with raw_wiki, +primary, and +heuristic. We fine-tune each model on SIB200. Hyperparameter settings are listed in Appendix A.
XLMR Continued Pretraining
Lang baseline raw_wiki +primary +heuristic

### v2

@@ 5.1 Monolingual Pretraining
To isolate the impact of data quality on downstream performance, we pretrain monolingual masked language models on 50 Wikipedias from the 4 tiers described in Section 4, selected based on the availability of downstream evaluation tasks. We employ three different versions of these Wikipedias as pretraining data: (1) the entire Wikipedia of the language without any filtering (raw_wiki), (2) filtered Wikipedia using the primary filters (+primary), and (3) further pruned using the heuristic filters (+heuristic), as described in §3. In addition to this, we include a control experiment (random) wherein we sample $n$ random articles from the Wikipedia, where $n$ is equal to the number of articles filtered in the primary filtering step. The full experimental pipeline is illustrated in Figure 6. All models are subsequently fine-tuned on MasakhaNER 2.0 Adelani et al. (2022), MasakhaPOS Dione et al. (2023), MasakhaNEWS Adelani et al. (2023), AfriSenti-Twitter Muhammad et al. (2023) and SIB200 Adelani et al. (2024), based on availability.
We employ small monolingual models so as to minimize the risk of confounding factors such as unbalanced multilingual data or tokenisation inequalities, while gaining training efficiency (Warstadt et al., 2023). We use the DeBERTa architecture He et al. (2021), largely due to its competitive performance in the low-data regime (Samuel et al., 2023). Details about our model and tokeniser hyperparameters, and training regime can be found in Appendix A.
Figure 7: Results for monolingual pretraining across 4 tasks for +primary and +heuristic. The y-axis plots the respective performance difference for each filtering method, with the gray dotted line corresponding to no difference. The x-axis plots the fraction of data removed (of articles $\cdot$ % of characters), to visualise the interaction between data filtering and performance. Points are coloured to denote tier categories - blue (Tier 1), orange (Tier 2), green (Tier 3), red (Tier 4).
Table 2 displays the average scores for all languages across filter methods and tasks. Here, we can observe that, while +primary generally yields similar performance as raw_wiki (MasakhaNER and AfriSenti-Twitter), it occasionally leads to better performance than no filtering at all (MasakhaNews). With the exception of AfriSenti-Twitter, both +primary and +heuristic perform better than random, indicating that the quality filters capture a notion of quality that is beneficial to model performance. In general, primary filtering appears to be more effective on average than heuristic filtering.
Task raw_wiki random +primary +heuristic
NER 0.81 0.77 0.81 0.80
News 0.80 0.76 0.82 0.80
SentiA 0.59 0.59 0.59 0.59
SIB200 0.63 0.63 0.62 0.50
Table 2: Average scores for all languages across filtering methods and tasks. Individual results are Appendix A
We can note from Figure 7 that for MasakhaNER and MasakhaNEWS, the removal of a large amount of data does not lead to a performance decline — in fact, yo actually enjoys a performance increase, despite 65% of the Wikipedia being removed. Interestingly, most languages that see a slight performance decrease happen to be allocated to Tier 1 (blue), and thus lose a smaller fraction of the data. We observe that +heuristic elicits a similar trend for MasakhaNEWS, but likewise note that the general performance decline for MasakhaNER, with respect to raw_wiki.
Interestingly, we find that all filtering strategies — including random— yield identical performance on AfriSenti-Twitter. We surmise that this could be due to a severe domain mismatch between the formal, encyclopaedic nature of Wikipedia and the social media-oriented nature of this particular dataset. As such, our assumptions about quality pretraining data do not directly apply here. We do note that ary is an exception, where the raw_wiki and random (see Table 10) distributions outperform quality filtering to a large degree. This can be explained by the fact that the ary subset in AfriSenti-Twitter contains script switching between Arabic and Latin. In our training setup, only raw_wiki and random models are able to tokenise Latin characters due to the script filtering. This connects to the discussion in §2.3 about what data quality actually means (more in §6)
Figure 8: Distribution of raw_wiki (blue), +primary (orange), +heuristic (green) and random (red) scores with monolingual pretraining across 50 languages in SIB200.
With SIB200 we are able to evaluate 50 languages, evenly spread across tiers. Thus, we can adequately compare the effectiveness of data quality filtering with respect to the existing quality of the Wikipedia. A KDE plot of the distribution of scores for each filtering method separated by tier can be found in Figure 8. Here, we observe that the distribution of the random scores is, on average, lower than the other three methods — with a more pronounced difference seen in Tiers 2, 3 and 4 than Tier 1. The underlying distribution of the Wikipedia, thus, has a meaningful effect on the impact of the quality filtering — as we also observed with MasakhaNEWS and MasakhaNER. We consider this to be a notable finding, since primary and heuristic filtering remove a larger proportion of articles for lower-ranked Wikipedias than those in Tier 1. This implies that the articles removed are of poor quality and are not beneficial to the training process. As a testament to this, we point to ceb in Figure 7, which has a 12% performance gain with +primary. We attribute this primarily to the deduplication step, which removes most bot generated articles.
5.2 Language adaptation
While the small-scale models of the previous section grant us full control of crucial variables relevant to our study, we are likewise interested how methods fare in a more common scenario. To this end, we turn to language adaptation: a common practice in the realm of low-resource and cross-lingual NLP concerned with adapting existing mono or multilingual models to languages unseen during pretraining Pfeiffer et al. (2020, 2022). Here, we follow the experimental protocol of Alabi et al. (2022) (see Appendix A) in adapting XLM-R base (Conneau et al., 2020) to unseen languages from our experimental cohort. We use the original XLM-R tokeniser and further pretrain XLM-R on individual Wikipedias across the raw_wiki, +primary, and +heuristic settings. Following this, we fine-tune each model on SIB200 and Belebele Bandarkar et al. (2024) — a cross-lingual reading comprehension bencmark with coverage over 122 languages. [^10] We fine-tune the models on the English fine-tuning data recommended in the Belebele paper, and evaluate it on the target languages in a zero shot manner. This setup is not possible with the monolingual models due to tokeniser restrictions.
Figure 9: Performance difference for XLMR+primary and XLMR+heuristic relative to XLMR+raw_wiki on SIB200 (with in-language fine-tuning) and Belebele (zero-shot).
Figure 10: Performance difference for +primary and +heuristic with multilingual pretraining relative to raw_wiki.
In general, we find that results for SIB200 in Figure 9 echo those observed for the monolingual models in Figure 7. In other words - while all languages (except mni) benefit from language adaptation (see Table 12), the effect of quality filtering is dependent upon the underlying data distribution. Notably, Tier 2, Tier 3 and Tier 4 benefit more from primary filtering, with ace achieving an 8% performance boost. Tier 1 Wikipedias, on the other hand, do not see much performance improvement from filtering on average. We make a special note of mni, which suffers a staggering 16% performance drop after language adaptation (see Table 12), raising many questions about the quality of that particular Wikipedia. This gap is notably reduced with primary filtering, however.
We observe the greatest benefit of quality filtering — especially primary filtering — in the Belebele evaluation, depicted in Figure 9. With the exception of bm, all languages experience a substantial performance boost, amounting to 5% on average. Previous work has shown that quality of pretraining data is more relevant to performance in a zero-shot setting than when fine-tuning the model on the target language (Tatariya et al., 2023). Thus, the higher quality +primary data can be expected to provide a benefit here. Beyond this, we observe that +heuristic filtering does not provide an advantage over primary filtering in either task. However, model performance in this setting is still very close to raw_wiki, despite the large-scale article loss of various Wikipedias.
5.3 Multilingual Pretraining
In our last downstream experiment, we aim to evaluate the effect of data quality filtering in a multilingual pretraining scenario. To do so, we train three multilingual masked language models on the top 100 Wikipedias in terms of article count, mimicking the setup of multilingual BERT Devlin et al. (2019). As with the monolingual and language adaptation experiments, we partition our data according to the raw_wiki, +primary and +heuristic settings. Across all models, all hyperparameters (see Appendix A) are kept constant, and each model is trained for the same amount of steps, allowing for comparability across settings. After training, we fine-tune these models on English task-specific data and evaluate zero-shot on Belebele and XNLI (Conneau et al., 2018).
Ultimately, we observe similar trends as before on Belebele (Figure 10), with the lower tiers showing more performance gains than Tier 1 with +primary. However, we do not observe the same degree of performance increase as depicted in Figure 9. For example, while ht enjoyed a 9% performance gain with XLMR in the +primary setting, it remains on par with raw_wiki after multilingual pretraining. On the other hand, other languages such as yo do not demonstrate any performance difference across settings. We interpret this to indicate that factors other than quality data can affect downstream performance for multilingual models, such as the language choice and composition. Going further, we note that +heuristic results also follow trends from earlier experiments, with most languages clustered around the dotted line — i.e. similar performance to raw_wiki. Notably, we find that war and ceb appear to benefit from the +heuristic step, despite retaining only on 10% and 18% of their data, respectively. With XNLI, we see that performance generally degrades after both quality filtering settings. However, it is important to note that most languages in XNLI have Wikipedias categorized under Tier 1. Notably, the performance for vi remains close to raw_wiki for both filtering methods, continuing to follow the overall trend.
6 Discussion and Recommendations
Effect of Quality Filtering
The quality filters tested in our experiments outperform random filtering, thus establishing that the primary and heuristic filters capture a notion of quality relevant to downstream performance. Performance after data quality filtering generally remains on par or better than no filtering at all. We look for a correlation between data pruning and accuracy with a Spearman’s correlation test between the amount of data removed in the primary filtering step, and the accuracy difference between raw_wiki and +primary. There is no significant correlation, which means that data quality pruning may not have a direct significant impact on performance. Findings from van Noord et al. (2024) and Artetxe et al. (2022) also echo this result. A substantial amount of bad data can be pruned without seeing any performance loss. In fact, Wikipedias in Tier 2, Tier 3 and Tier 4 benefit more from primary filtering than those in Tier 1. Thus, we recommend data quality filtering as a means for resource efficient training, in addition to the benefit of better performance for many languages.
Heuristic filtering, on the other hand, yields mixed results. This cannot be entirely attributed to aggressive data filtering, as Tier 1 Wikipedias have more performance decline than those in lower tiers. We speculate that this discrepancy arises because heuristic metrics are computed at the article level, whereas model training occurs at the chunk level. Likewise, the heuristic filters that we deploy can be considered rather generic in nature, accounting only for distributionally-conditioned word and character-level statistics. It is likely that, for language-specific applications, one may include specialized heuristics that capture more information about the data domain — e.g. specific keywords or phrases — and choose specific thresholds for these purposes. Furthermore, +primary may already possess a well-balanced data distribution, making further filtering detrimental. Regardless, we would like to note that heuristic filtering still has its merits for many languages, thus warranting further analysis in future work.
Between monolingual and multilingual pretraining, the advantages of quality filtering are seen most in the monolingual setting, especially in the zero-shot scenario. Trends noted above remain observable in the multilingual models but less pronounced, likely due to cross-lingual interactions influencing the outcomes. We recommend always using data quality filters when fine-tuning data is not available, or when adapting existing models to a low-resource language, especially if the Wikipedia is in Tier 2, Tier 3 and Tier 4.
Quality versus Size
Figures 7, 9 and 10 demonstrate that primary and heuristic filtering steps can potentially prune up to 60-80% of the data in a bad Wikipedia. This is also observed for lower resource languages with small Wikipedias (see yo, nso, lg, pag, to name a few) where the assumption might be that there is already very little data available, rendering any further filtering to be detrimental. We demonstrate that even for such small Wikipedias, removal of bad data is desirable, even beneficial to performance on downstream tasks. Moreover, when compared with the XLM-R model adapted to the same data (for e.g., monolingual raw_wiki versus XLM-R+raw_wiki), 23 monolingual models either match or outperform XLM-R on SIB200 — all with a fraction of the data size and capacity. Moreover, our results for AfriSenti-Twitter are competitive with those reported by Muhammad et al. (2023) on larger multilingual models. This indicates that a smaller concentration of monolingual data trained with fewer parameters can yield better performance than a large multilingual corpus that includes the same data. Therefore, low-resource languages benefit from smaller, more efficient models trained on a fraction of the compute with quality filtering being paramount to achieving good performance and getting the most out of limited data.
Quality versus Language
Although the results discussed in this study largely remain consistent in aggregate, zooming in on individual languages reveals the large benefits of the filtering process for some languages. For example, om and ti benefit from a substantial performance boost on MasakhaNEWS after primary filtering. Although +heuristic generally does not prove to be as competitive as +primary, certain languages — such as war and jv with multilingual pretraining — still benefit considerably from the process. As such, we emphasize that there there exists no one-size-fits-all data filtering approach. Instead, we advocate for practitioners to move away from seeking general definitions of data quality and focus rather on the examination of filtering strategies for individual Wikipedias. Concretely, data quality decisions with regard to user-generated multilingual corpora such as Wikipedia should be made on a per-language basis, despite the theoretical domain to which they belong. Furthermore, this approach can benefit from leveraging typological features and databases to create more language-specific data filtering pipelines, which is a promising direction for future work.
Quality and Downstream Tasks
Aside from a pretraining resource, Wikipedia is an equally important primary source for multilingual, task-specific dataset creation. While we have focused largely on the former, we highlight some implications of data quality filtering for the latter.
Lent et al. (2024) examined the quality of Creole Wikipedias, observing that many articles contain short and templatic entries as a result of machine generation. In this work, we have engaged in discussions regarding the benefit of removing such articles in a pretraining setting. However, Lent et al. (2024) note that the templatic nature of such Wikipedias, in fact, happen to facilitate their annotation process in creating a relation-classification dataset. In such scenarios, our filtering pipeline can be employed to describe and identify useful articles. For example, in this case Lent et al. (2024)’s experiments, heuristic filtering can be employed to identify articles that are very short, with low entropy, and thus are more likely to be templatic and generated from a knowledge base.
In contrast to this, Kreutzer et al. (2022a) discuss the pervasive impact of low data quality in downstream evaluation with the example of machine translation. This is particularly salient in the low-resource regime, where Wikipedia is often the only source of data — potentially leading to flawed evaluation datasets and misleading conclusions. Lignos et al. (2023) highlight quality issues with the WikiAnn dataset Pan et al. (2017), reporting much noise present in the data, along with many examples that either are very dense in mentions or consist entirely of single entity mentions, particularly for African languages. This could likely stem from the broader quality issues endemic to Wikipedia.
7 Conclusion
In this study, we challenged the assumption of Wikipedia as a high quality resource in multilingual NLP, and provided a guide to utilizing it as a pretraining resource in NLP. We show that much of Wikipedia contains numerous quality issues such as duplicate articles, bot-generated content, and foreign scripts, among other issues, which can be filtered out as a means to achieve good performance with less compute. For select languages, this process can even lead to performance improvement. Overall, we advocate towards a shift in perspective from seeking general definitions of data quality to approaching quality issues from a domain or language specific lens.
8 Acknowledgements
Appendix A Training Hyperparameters and Finetuning Results
Hyperparameters for the pretraining and fine-tuning of monolingual models are in Table 3. We use the SentencePiece package Kudo and Richardson (2018) to train monolingual unigram tokenisers for each language. We train two tokenisers per language: one on the raw_wiki with no filtering, which is used to train the raw_wiki models; the second one on the data remaining after script filtering and deduplication, which we use to train all other models. We use an approximation with Zipf’s law Zipf (1949) on each Wikipedia to estimate the vocabulary size for each model, mentioned in Table 6. Hyperparameters for XLM-R language adaptation are in Table 4. For both cases, we checkpoint and use the best model on a left-out validation set to fine-tune after pretraining. For multilingual pretraining, we train 2 WordPiece tokenizers on raw_wiki and +primary (used to train both +primary and +heuristic models) with a vocabulary size of 110k. Languages are listed in Table 7 and hyperparameters in Table 5.
Tiny DeBERTa Hyperparameters
Pretraining
Model Type deberta
Hidden Size 312
Intermediate Size 1200
Attention Heads 12
Hidden Layers 4
Hidden Act gelu
Max Length 128
Task MLM
Mask Probability 0.4
Padding Strategy longest
Grad. Accumulation Steps 4
Batch size 8 to 32
Learning Rate 1.00e-3
Epochs 100
Finetuning
Max Length 128
Epochs 20
Batch Size 8
Learning Rate 5.00e-5
Padding Strategy max_length
Table 3: Hyperparameters for monolingual pretraining and fine-tuning. While we used 100 epochs as a default, for Wikipedias that are too large (nl, ro, id, hu, uk, vi, ceb, war, tt, azb, min) we pretrain only for 2 epochs. Similarly, while we used a batch size of 8 for all models, for the larger Wikipedias we use a batch size of 32. All models are trained on 1 GPU.
XLMR Hyperparameters
Language Adaptation
Model Type xlm-roberta
Task MLM
Max Length 512
Mask Probability 0.15
Epochs 3
Batch Size 8
Learning Rate 5.00e-5
Padding strategy Longest
Grad. Accumulation Steps 1
Finetuning - SIB200
Max Length 512
Epochs 20
Batch Size 32
Learning Rate 5.00e-5
Padding Strategy max_length
Finetuning - Belebele
Max Length 128
Epochs 3
Batch Size 8
Learning Rate 5.00e-5
Padding Strategy max_length
Table 4: Hyperparameters for XLM-R language adaptation (based on Alabi et al. (2022)) and fine-tuning.
Multilingual Pretraining
Pretraining
Sampling Temperature 1.43
Model Type bert
Task MLM
Max Length 128
Mask Probability 0.12
Steps 200,000
Batch Size (per device) 256
Learning Rate 1.00e-3
Padding strategy Longest
Grad. Accumulation Steps 4
Finetuning - Belebele
Max Length 128
Epochs 3
Batch Size 8
Learning Rate 5.00e-5
Padding Strategy max_length
Finetuning - MNLI
Max Length 128
Epochs 3
Batch Size 32
Learning Rate 2.00e-5
Padding Strategy max_length
Table 5: Hyperparameters for multilingual pretraining and fine-tuning, based on mBERT and Izsak et al. (2021). All three models are trained on 4 H100 GPUs.
Estimated Vocabulary Size
Lang Vocab Wiki-ID Vocab
ace 3505 mr 14708
am 5234 nl 60011
ary 4502 nso 2801
awa 2190 ny 2472
azb 13996 om 3436
ban 6624 pag 2170
bm 1485 pcm 2595
ceb 73124 rn 1446
ckb 11451 ro 36329
ee 1770 rw 5725
fon 1426 shn 5784
ha 13126 sn 5227
ht 10836 so 6117
hu 46205 sw 12529
id 41725 te 23175
ig 11609 ti 1150
kk 22496 tn 3783
ks 2248 ts 1817
lg 4528 tt 26047
lij 5462 tw 4745
ln 2509 uk 62626
mai 4793 vi 43806
mi 3451 war 25803
min 14783 wo 3384
mni 3351 yo 6128
zu 4439
Table 6: Estimated vocabulary size used to train monolingual tokenisers for each Wikipedia in our experimental pipeline.
Languages Included in Multilingual Pretraining
Language Wiki-ID Language Wiki-ID
Afrikaans af Korean ko
Albanian sq Latin la
Arabic ar Latvian lv
Aragonese an Lithuanian lt
Armenian hy Lombard lmo
Asturian ast Low German nds-nl
Azerbaijani az Luxembourgish lb
Bashkir ba Macedonian mk
Basque eu Malagasy mg
Bavarian bar Malay ms
Belarusian be Malayalam ml
Bangla bn Marathi mr
Bishnupriya bpy Minangkabau min
Bosnian bs Nepali ne
Breton br Newari new
Bulgarian bg Norwegian Nynorsk nn
Burmese my Occitan oc
Catalan ca Persian fa
Chechen ce Piedmontese pms
Chinese zh Polish pl
Cantonese zh-yue Portuguese pt
Cebuano ceb Punjabi pa
Chuvash cv Romanian ro
Croatian hr Rusyn rue
Czech cs Scots sco
Danish da Serbo-Croatian sh
Dutch nl Sicilian scn
English en Slovak sk
Estonian et Slovenian sl
Finnish fi South Azerbaijani azb
French fr Spanish es
Galician gl Sundanese su
Georgian ka Swahili sw
German de Swedish sv
Greek el Filipino tl
Gujarati gu Tajik tg
Haitian Creole ht Tamil ta
Hebrew he Tatar tt
Hindi hi Turkish tr
Hungarian hu Ukrainian uk
Icelandic is Uzbek uz
Ido io Vietnamese vi
Indonesian id Volapük vo
Irish ga Waray war
Italian it Welsh cy
Japanese ja Western Frisian fy
Javanese jv Western Panjabi pnb
Kannada kn Yoruba yo
Kazakh kk Thai th
Kyrgyz ky Mongolian mn
Table 7: Wikipedias included in the pretraining of the multilingual models.
@@ Appendix A Training Hyperparameters and Finetuning Results
Lang raw_wiki +primary +heuristic random
pcm 0.77±0.00 0.76±0.00 0.76±0.00 0.74±0.00
ha 0.82±0.00 0.81±0.00 0.81±0.00 0.75±0.01
ig 0.85±0.00 0.85±0.00 0.84±0.00 0.81±0.01
sn 0.92±0.00 0.92±0.00 0.92±0.00 0.90±0.00
rw 0.76±0.05 0.79±0.01 0.77±0.05 0.74±0.01
tw 0.76±0.00 0.76±0.01 0.74±0.00 0.71±0.00
wo 0.75±0.01 0.73±0.01 0.72±0.01 0.73±0.01
ny 0.87±0.00 0.87±0.00 0.86±0.00 0.85±0.00
ee 0.81±0.00 0.81±0.01 0.79±0.00 0.78±0.01
bm 0.74±0.01 0.74±0.01 0.72±0.01 0.73±0.00
fon 0.74±0.01 0.74±0.00 0.76±0.01 0.71±0.01
sw 0.90±0.00 0.90±0.00 0.90±0.00 0.88±0.00
zu 0.83±0.01 0.82±0.01 0.81±0.01 0.77±0.01
lg 0.84±0.00 0.85±0.00 0.83±0.00 0.80±0.01
tn 0.80±0.01 0.80±0.00 0.79±0.01 0.63±0.01
yo 0.84±0.00 0.84±0.00 0.83±0.00 0.80±0.00
Avg 0.81 0.81 0.8 0.77
@@
wiki raw_wiki +primary +heuristic random
pcm 0.90±0.04 0.91±0.00 0.87±0.04 0.88±0.02
ha 0.88±0.01 0.86±0.01 0.86±0.01 0.84±0.01
ig 0.85±0.01 0.85±0.01 0.84±0.01 0.81±0.01
sn 0.89±0.01 0.90±0.01 0.89±0.01 0.87±0.01
so 0.71±0.01 0.71±0.01 0.70±0.01 0.59±0.03
ln 0.78±0.01 0.76±0.01 0.76±0.01 0.74±0.02
om 0.67±0.30 0.82±0.01 0.81±0.30 0.71±0.04
ti 0.60±0.04 0.65±0.02 0.62±0.04 0.56±0.02
sw 0.81±0.01 0.82±0.00 0.82±0.01 0.75±0.01
am 0.84±0.04 0.87±0.01 0.85±0.04 0.83±0.01
lg 0.86±0.01 0.86±0.02 0.84±0.01 0.84±0.02
rn 0.75±0.01 0.76±0.02 0.71±0.01 0.71±0.01
yo 0.84±0.03 0.86±0.01 0.86±0.03 0.82±0.01
Avg 0.8 0.82 0.8 0.76
@@
wiki raw_wiki +primary +heuristic random
pcm 0.61±0.01 0.61±0.01 0.61±0.01 0.60±0.01
ha 0.72±0.01 0.72±0.01 0.73±0.01 0.72±0.01
ig 0.74±0.00 0.75±0.00 0.75±0.00 0.74±0.01
rw 0.60±0.04 0.61±0.01 0.59±0.04 0.60±0.01
tw 0.62±0.01 0.64±0.01 0.64±0.01 0.63±0.01
ts 0.48±0.04 0.49±0.03 0.50±0.04 0.50±0.01
sw 0.55±0.03 0.59±0.02 0.56±0.03 0.53±0.04
am 0.40±0.07 0.41±0.04 0.46±0.07 0.45±0.10
ary 0.46±0.01 0.40±0.01 0.40±0.01 0.44±0.01
yo 0.68±0.01 0.68±0.01 0.68±0.01 0.64±0.01
Avg 0.59 0.59 0.59 0.59
@@
wiki raw_wiki +primary +heuristic random
@@
XLMR Continued Pretraining (SIB200)
wiki baseline raw_wiki +primary +heuristic